# Learning a Shape-adaptive Assist-as-needed Rehabilitation Policy from Therapist-informed Input

### v2

@@ Abstract
Therapist-in-the-loop robotic rehabilitation has shown great promise in enhancing rehabilitation outcomes by integrating the strengths of therapists and robotic systems. However, its broader adoption remains limited due to insufficient safe interaction and limited adaptation capability. This article proposes a novel telerobotics-mediated framework that enables therapists to intuitively and safely deliver assist-as-needed (AAN) therapy based on two primary contributions. First, our framework encodes the therapist-informed corrective force into via-points in a latent space, allowing the therapist to provide only minimal assistance while encouraging patient maintaining own motion preferences. Second, a shape-adaptive ANN rehabilitation policy is learned to partially and progressively deform the reference trajectory for movement therapy based on encoded patient motion preferences and therapist-informed via-points. The effectiveness of the proposed shape-adaptive AAN strategy was validated on a telerobotic rehabilitation system using two representative tasks. The results demonstrate its practicality for remote AAN therapy and its superiority over two state-of-the-art methods in reducing corrective force and improving movement smoothness.
@@ I Introduction
Rehabilitation robots can provide precise and intensive therapies for patients with neurological injuries, offering a promising solution to the shortage of healthcare resources [1]. Numerous robotic systems [2, 3] and control methodologies [4],[5] have been developed to promote robot-assisted rehabilitation. Particularly, adaptive control and machine learning techniques have shown promising results in programming flexible therapy strategies tailored to patients’ individual needs, known as assist-as-needed (AAN) strategies [6, 7]. Nevertheless, the flexibility of existing AAN strategies cannot match the capability of a skilled therapist due to the difficulty of acquiring sufficient knowledge and programming experience-based adjustments [8],[9],[10]. To address this limitation, therapist-in-the-loop robotic rehabilitation has emerged as a paradigm that integrates therapists’ expertise into robot controllers [11],[12],[13]. However, most existing robotic systems still struggle to complete effective therapist-in-the-loop rehabilitation training because of their limited interaction and adaptation capabilities.
@@
Firstly, current robotic rehabilitation systems lack effective mutual interaction to safely integrate therapists’ expertise into robot execution [14],[15]. Robot controllers are typically designed to follow the predefined reference based on the measured interaction forces and motion tracking for desired rehabilitation targets [8]. Therapists, in turn, contribute by setting rehabilitation goals and modulating the therapy strategies tailored to individual patients for robot execution based on previous expertise in evaluating individual patient status [10]. In practice, therapists typically demonstrate therapeutic movements/forces for robot-mediated rehabilitation using the same robot in a teach-and-play mode [16]. By contrast, telerobotics-mediated systems enable therapist and patient to operate separate devices, enabling real-time mutual interaction [17],[12]. For instance, therapists utilize a haptic device to perceive the interaction force between the patient and the robot. Existing robotic telerehabilitation systems employ additional sensors and estimation methods to achieve concurrent demonstration and therapy [18]. However, no current telerobotic-mediated systems can intuitively incorporate therapist’s input into robot control without relying on reprogramming.
@@
Secondly, current robotic rehabilitation strategies struggle to deliver effective AAN therapy due to their limited adaptability. The key challenge in developing AAN training strategies is to encourage active participation while only providing the necessary level of assistance [8, 19, 5]. Patient’s voluntary output forces and motion tracking errors are commonly used as metrics to assess patient engagement and guide the design of robot controllers for various rehabilitation goals [20, 6, 21]. While therapists‘ expertise allows for comprehensive evaluation of the patients’ status through multi-modal feedback, therapist’s corrective actions are essential to promote therapist-level adaptive assistance [22]. Numerous studies have leveraged probabilistic models to reproduce therapist expertise for reference trajectory shaping and force generation in AAN therapy [4]. In these methods, rehabilitation outcomes are defined by the therapist and the corresponding references are generated for robot controllers [23, 5]. However, the effectiveness of therapy remains constrained by the difficulty of balancing robotic assistance with patients’ active participation. Therefore, a promising therapist-in-the-loop AAN training strategy that maximizes patient engagement while seamlessly integrating therapist-informed corrections remains necessary.
@@
This article proposes a novel therapist-in-the-loop rehabilitation framework that leverages two isomorphic collaborative robots to deliver the AAN movement therapy. The contributions of this study are twofold. First, therapist-informed corrective forces are encoded as via-points in a latent space, ensuring that the therapist only provides the minimal necessary assistance instead of re-planning the entire reference trajectory or force profile. Second, a shape-adaptive AAN rehabilitation policy is learned to partially and progressively deform the reference trajectory based on the encoded patient motion preferences and therapist-informed via-points.
@@ II-A AAN Training Strategies
AAN training strategies have been investigated at multiple levels for patients with various rehabilitation goals by modulating the amount of assistance based on the biomechanical status of the patient or task completion performance [20],[24]. At the force level, adaptive controllers have been developed to successfully modify the level of assistance in real time [6],[25]. Furthermore, multi-modal adaptive controllers have been proposed to provide flexible assistive force across multiple training modes [26],[27]. Beyond the adaptation of assistive force, impedance control, a popular compliant control approach, was utilized to regulate the patient-robot interaction dynamics during movement therapy [23]. Variable impedance controllers (VICs) have been commonly developed to enable impedance-level AAN therapy, where individual reference stiffness parameters are predefined according to patient’s task completion performance [28],[23],[29]. These training strategies typically achieved the force-level and impedance-level AAN therapy given a fixed reference trajectory [28]. By contrast, the individualized reference trajectory deformation for motion-level AAN therapy has demonstrated the improved stability [30]. The reference trajectory was modified based on optimization deformation for patients with different mobility disabilities and different recovery stages [31]. Additionally, probabilistic model learning methods, Gaussian Mixture Model (GMM)/Gaussian Mixture Regression (GMR) [11], were applied to reproduce the therapist-demonstrated reference trajectories [8]. More recently, moving beyond single rehabilitation tasks, task-level AAN strategies have been explored. These approaches employ linear Gaussian policies or neural network–based policies to encode mappings from task contexts to a latent space of lower-level policies for robot execution [32, 33].
Most of existing AAN rehabilitation strategies rely on predefined metrics and corresponding robot controllers for execution. However, these approaches cannot match the flexibility of skilled therapists. An ideal AAN training strategy should combine the strengths of therapists’ expertise with the precise execution of robots. Crucially, the therapist should provide only the minimal necessary guidance, thereby maximizing patients’ active participation.
@@ II-B Therapist-in-the-loop Rehabilitation Training
Therapist-in-the-loop training strategies rely on the therapist to demonstrate therapeutic movements or force profiles for robot execution [8, 12]. The ability to modify therapy movements or intensity has been validated in the teach-and-play mode [4]. In the demonstration phase, the therapeutic movement or force delivered by therapists is recorded and fitted by probabilistic models [11],[12]. In the reproduction phase, the learned probabilistic model can reproduce the reference trajectory or force profiles without the therapist in the loop [16]. The therapeutic intensity was modulated by modifying controller parameters of the patient-side robot to provide AAN therapy [5]. Additionally, leader-follower robotic systems have been developed for telerobotic-mediated rehabilitation to provide concurrent demonstration and therapy. Therapists use a haptic device to demonstrate the reference trajectory and perceive patient-side interaction force [10], which allows for adjusting the therapeutic movement and force in real-time to ensure effective therapy [10],[34]. External sensors, such as EMG sensors, are amounted to infer control parameters for modulating the therapy intensity.
In summary, compared with methods using a single robot, a telerobotics-mediated rehabilitation system enables the decoupling of evaluation and therapeutic treatment. Unfortunately, in existing strategies, therapists are required to modulate entire therapeutic movements or force profiles [12]. By contrast, an ideal AAN therapy strategy would encode therapist-informed corrective force into a latent space to smoothly guide patient movement therapy.
III Problem Formulation
@@
The objective of the proposed controller is to enable the therapist to provide necessary guidance for completing the AAN therapy with patients. As illustrated in Fig. 1, the telerobotic system consists of two identical collaborative robots to complete the therapist-in-the-loop training. One robot, located at the patient side (patient-side robot), is driven using an impedance controller to guide the patient to compliantly complete the movement therapy. Another robot, located on the therapist side (therapist-side robot), is driven using another impedance controller to allow him/her to perceive the actual trajectories of the patient-side robot. Two GUIs are developed for the mutual visual interaction between therapists and patients. For effective AAN therapy, the reference trajectory is adapted to encourage the active participation of patients where the patient motion preferences are considered and maintained. Furthermore, the therapist-side robot isolates the therapist-informed corrective force from the patient-side interaction force, which enables it to provide accurate guidance.
@@
This article focuses on movement therapy given a set of predefined desired motions (see Fig. 1), following the implementation in Algorithm 1. Once a rehabilitation task is selected by the therapist, the desired motion $\bm{x}_{e}$ is displayed on GUIs to provide task instructions for both therapists and patients. The robot-assisted rehabilitation relies on repetitive implementations based on two lower-level interactive controllers depicted in Section IV-A. During each therapy iteration, the patient’s actual trajectories $\bm{x}_{t}$ are measured from the patient-side robot. The patient motion preferences are encoded from recent actual trajectories by learning a probabilistic model (see Section IV-B). To minimize therapist assistance and encourage patient active participation, therapist-informed via-points are extracted from the measured corrective force instead of transmitting the whole force profile or reference trajectory for patient-side robot control. The corresponding reference trajectory $\bm{x}_{r}$ for subsequent therapy iteration is then partially deformed according to the therapist-informed via-points, while preserving the patient’s motion preferences in the remaining segments (see Section IV-C). Most importantly, the therapist’s input therapeutic skills are encoded by using via-points a latent space according to the feature of patient motion preferences. Finally, a partial least squares regression function is fitted to reproduce the therapist’s input therapeutic skills (see Section IV-D).
@@ IV-A Lower-level Interactive Control
The robot dynamics in $n$ Degree of Freedom (DoF) Cartesian space is written as,
@@
where $\bm{M}_{C,i}\in\mathbb{R}^{n\times n}$ denotes the inertia matrix of the robot, and $\bm{c}_{C,i}\in\mathbb{R}^{n}$ and $\bm{g}_{C,i}\in\mathbb{R}^{n}$ represent the Coriolis/centrifugal and the gravity vector, respectively. $\bm{f}_{cmp,i}\in\mathbb{R}^{n}$ is the control force which compensates the Coriolis/centrifugal and the gravity vectors. $\bm{f}_{ctrl,i}\in\mathbb{R}^{n}$ is the control command. Moreover, $\bm{f}_{ext,i}\in\mathbb{R}^{n}$ is the force human user applies on the robot. The subscript $_{i}$ is indexed by $i\in\{P,Th\}$, where $_{P}$ indicates the robot at the patient side, and $_{Th}$ for the therapist-side robot.
@@ IV-A1 Patient-side Robot Control
On the patient side, the robot is following the reference trajectory generated by the therapist input, thus the controller at the patient side can be written as
@@ IV-A2 Therapist-side Robot Control
The robot on the therapist side is following the patient-side robot such that the therapist understand the situation of the patient. Thus the controller on the therapist-side robot is designed as,
@@
where the $\bm{K}_{Th},\bm{D}_{Th}\in\mathbb{R}^{n\times n}$ are the positive definite stiffness and damping matrices for the therapist-side robot, respectively. and the force interaction force between the therapist $\bm{f}_{ext,Th}$ will be the input of the trajectory generation.
@@ IV-B Patient Motion Preferences Encoding
For each therapy iteration, the impedance control in (3) is utilized to complete the movement therapy following a reference trajectory $\bm{x}_{r}$ for $J$ times. Most recent actual trajectories of patient (e.g. gray lines in the Fig. 2(a)) are collected and down-sampled into a set of waypoints with length NN for learning a propabilistic model to encode the motion preferences. The dataset $\mathcal{D}_{P}$ is constructed as $\mathcal{D}_{P}=\{\{\bm{\xi}_{n,j}^{i},\bm{\xi}_{n,j}^{o}\}_{n=1}^{N}\}_{j=1}^{J}$. $\bm{\xi}_{n,j}^{i}\in\mathbb{R}^{d_{i}}$ and $\bm{\xi}_{n,j}^{o}\in\mathbb{R}^{d_{o}}$ represent the input and output vectors with dimensions of $d_{i}$ and $d_{o}$. This study focuses on time-driven trajectories for movement therapy, where the input $\bm{\xi}_{n,j}^{i}$ corresponds to time $\bm{t}$, and the output $\bm{\xi}_{n,j}^{o}$ corresponds to the end-effector position $\bm{x}$ of patient-side robot. GMMs are used to encode the mean and variance of motion preferences using collected waypoints in $\mathcal{D}^{i}_{P}$. The covariance encodes the variability of the actual trajectories. The resulting joint probability distribution of time $\bm{t}$ and position $\bm{x}$ is given as
$\centering\small\mathcal{P}(\bm{t},\bm{x})=\sum\nolimits_{c=1}^{C}\varpi_{c}\mathcal{N}(\bm{t},\bm{x}|\bm{\mu}_{c},\bm{\Sigma}_{c})$ (4)
where $C$ is the number of Gaussian components, which controls the smoothness of the fitted mean trajectory. The Gaussian component parameters $\{\bm{\mu}_{c},\bm{\Sigma}_{c},\varpi_{c}\}_{c=1}^{C}$ are characterized by its mean $\bm{\mu}_{c}$, covariance $\bm{\Sigma}_{c}$, and weight $\varpi_{c}$, which can be estimated from $\mathcal{D}_{P}$ using the Expectation-Maximization algorithm [35] (see green line and ellipses in Fig. 2(a)). The conditional probabilistic trajectory, denoted by $\{\widehat{\bm{x}}_{n}\}_{n=1}^{N}$, is then retrieved via GMR as
$\centering\small\mathcal{P}(\widehat{\bm{x}}_{n}|\bm{t}_{n})\sim\mathcal{N}(\widehat{\bm{\mu}}_{n},\widehat{\bm{\Sigma}}_{n})$ (5)
where $\widehat{\bm{\mu}}_{n}$ and $\widehat{\bm{\Sigma}}_{n}$ are the conditional mean and covariance that can be calculated using the estimated GMMs parameters.

### v3

@@ Abstract
Therapist-in-the-loop robotic rehabilitation has shown great promise in enhancing rehabilitation outcomes by integrating the strengths of therapists and robotic systems. However, its broader adoption remains limited due to insufficient interaction and limited adaptation capability. This article proposes a novel telerobotics-mediated framework that enables therapists to intuitively deliver assist-as-needed (AAN) therapy based on two primary contributions. First, our framework encodes the therapist-informed corrective force into via-points in a latent space, allowing therapists to provide only minimal assistance while encouraging patients maintaining their own motion preferences. Second, a shape-adaptive AAN rehabilitation policy is learned to partially and progressively deform the reference trajectory for movement therapy based on encoded patient motion preferences and therapist-informed via-points. The effectiveness of the proposed shape-adaptive AAN strategy was validated on a telerobotic rehabilitation system using two representative tasks. The results demonstrate its practicality for remote AAN therapy and its superiority over two state-of-the-art methods in reducing corrective force and improving movement smoothness.
@@ I Introduction
Rehabilitation robots can provide precise and intensive therapies for patients with neurological injuries, offering a promising solution to the shortage of healthcare resources [1]. Numerous robotic systems [2, 3] and control methodologies [4],[5] have been developed to promote robot-assisted rehabilitation. Particularly, adaptive control and machine learning techniques have shown promising results in programming flexible therapy strategies tailored to patients’ individual needs, known as assist-as-needed (AAN) strategies [6, 7]. Nevertheless, the flexibility of existing AAN strategies cannot match the capability of a skilled therapist due to the difficulty of acquiring sufficient knowledge and programming experience-based adjustments [8],[9]. To address this limitation, therapist-in-the-loop robotic rehabilitation has emerged as a paradigm that integrates therapists’ expertise into robot controllers [10],[11]. However, most existing robotic systems still struggle to perform effective therapist-in-the-loop rehabilitation training because of their limited interaction and adaptation capabilities.
@@
First, current robotic rehabilitation systems lack effective mutual interaction to safely integrate therapists’ expertise into robot execution [12],[13]. Robot controllers are typically designed to track the predefined reference trajectories based on the measured interaction forces and actual trajectories for desired rehabilitation goals [8]. Therapists, in turn, contribute by setting rehabilitation goals and modulating the therapy strategies tailored to individual patients for robot execution based on previous expertise in evaluating patient status [9]. Therapists typically demonstrate therapeutic movements/forces for robot-mediated rehabilitation using the same robot in a teach-and-play mode [14]. By contrast, telerobotics-mediated systems enable therapist and patient to operate separate devices, enabling real-time mutual interaction [15],[11]. Therapists can utilize a haptic device to perceive the interaction force between the patient and the robot [8]. Additional sensors and estimation methods were employed to achieve concurrent demonstration and therapy [16]. However, no current telerobotic-mediated systems can intuitively incorporate therapist’s input into robot control without relying on reprogramming.
@@
Second, current robotic rehabilitation strategies struggle to deliver effective AAN therapy due to their limited adaptability. The key idea of developing AAN training strategies is to encourage active participation by only providing the necessary level of assistance [8, 17, 5]. Patients’ voluntary output forces and motion tracking errors are commonly used as metrics to assess patient engagement and guide the design of robot controllers for various rehabilitation goals [18, 6, 19]. While therapists’ expertise allows for comprehensive evaluation of the patients’ status through multi-modal feedback, therapists’ corrective actions are essential to promote therapist-level adaptive assistance [20]. Numerous studies have leveraged probabilistic models to reproduce therapist expertise for shaping reference trajectories and force profiles in AAN therapy [4]. In these methods, rehabilitation goals are defined by the therapist and the corresponding references are generated for robot controllers [7, 5]. However, the effectiveness of therapy remains constrained by the difficulty of balancing robotic assistance with patients’ active participation. Therefore, a promising therapist-in-the-loop AAN training strategy that maximizes patient engagement while seamlessly integrating therapist-informed corrections remains necessary.
@@
This article proposes a novel therapist-in-the-loop rehabilitation framework that leverages two isomorphic collaborative robots to deliver AAN therapy. The contributions are twofold. First, therapist-informed corrective forces are encoded as via-points in a latent space, enabling the therapist to provide only the minimal necessary assistance instead of re-planning the entire reference trajectory or force profile. Second, a shape-adaptive AAN rehabilitation policy is learned to partially and progressively deform the reference trajectory based on the encoded patient motion preferences and therapist-informed via-points.
@@ II-A AAN Training Strategies
AAN training strategies have been investigated at multiple levels for patients with various rehabilitation goals by modulating the amount of assistance according to the biomechanical status of the patient or task performance [18],[21]. At the force level, adaptive controllers have been developed to successfully modify the level of assistance in real time [6],[22]. Furthermore, multi-modal adaptive controllers have been proposed to provide flexible assistive force across multiple training modes [23]. Beyond adjusting the assistive force, impedance control, a compliant control method, was utilized to regulate the patient-robot interaction dynamics during movement therapy [7]. Variable impedance controllers have been commonly developed to enable impedance-level AAN therapy, where individual reference stiffness parameters are predefined according to patient’s task performance [7],[24],[25]. These methods mainly achieved the force-level and impedance-level AAN therapy given a fixed reference trajectory. By contrast, the individualized reference trajectory deformation for motion-level AAN therapy has been proven to improve the stability [26]. The reference trajectory was modified based on optimization-based deformation for patients with different mobility disabilities and at different recovery stages [27]. Additionally, probabilistic learning methods, Gaussian Mixture Model (GMM) and Gaussian Mixture Regression (GMR) [10], were applied to reproduce the therapist-demonstrated reference trajectories [8]. More recently, moving beyond single rehabilitation tasks, task-level AAN strategies have been explored. Linear Gaussian policies [28] or neural network–based policies [29] were developed to encode mappings from task contexts to a latent space of lower-level robot policies, which are formulated based on Kernelized Movement Primitive (KMP) [30]. Unfortunately, most existing approaches cannot match the flexibility of skilled therapists. An ideal AAN training strategy should combine the strengths of therapists’ expertise with the precise execution of robots. Crucially, the therapist should provide only the minimal necessary guidance, thereby maximizing patients’ active participation.
@@ II-B Therapist-in-the-loop Rehabilitation Training
Therapist-in-the-loop training strategies rely on demonstrating or modifying force profiles, therapeutic movements and therapeutic intensity following a teach-and-play paradigm [8, 11]. During the demonstration phase, the therapeutic movements or force profiles delivered by therapists are recorded and fitted by probabilistic models [10],[11]. In the reproduction phase, the learned probabilistic model can generate reference trajectories or force profiles without the therapist in the loop [14]. The therapeutic intensity is modulated by modifying controller parameters of the patient-side robot to perform AAN therapy [5]. Additionally, leader-follower robotic systems have been developed for telerobotic-mediated rehabilitation to enable concurrent demonstration and therapy. Therapists use a haptic device to demonstrate reference trajectories and to perceive patient-side interaction forces [9], allowing for adjusting therapeutic movements and force profiles in real-time to ensure effective therapy [8]. EMG sensors were employed to infer controller parameters for modulating therapy intensity. The telerobotics-mediated rehabilitation system enables the decoupling of evaluation and therapeutic treatment. Unfortunately, in existing strategies, therapists are typically required to modulate entire therapeutic movements or force profiles [11]. By contrast, an ideal AAN therapy strategy would encode therapist-informed corrective forces to guide the movement therapy.
III Overview Framework
@@
The objective of the proposed framework is to enable therapists to provide the necessary guidance for delivering the AAN therapy to patients. As illustrated in Fig. 1, the telerobotic system consists of two identical collaborative robots to perform the therapist-in-the-loop training. One robot, located at the patient side (patient-side robot), is driven by an impedance controller to guide the patient during movement therapy. The other robot, located on the therapist side (therapist-side robot), is driven by another impedance controller, allowing him/her to perceive the actual trajectories of the patient-side robot. Two graphical user interfaces (GUIs) are developed to enable mutual visual interaction between therapists and patients. For AAN therapy, the reference trajectory is adapted to encourage active participation of patients where the patient motion preferences are maintained. Furthermore, the therapist-side robot isolates the therapist-informed corrective force from the patient-side interaction force, enabling it to provide accurate guidance.
@@
This article focuses on movement therapy given a set of predefined desired motions [25]. Once a rehabilitation task is selected by the therapist (see Fig. 1), the desired motion $\bm{x}_{e}$ is displayed by GUIs to provide task instructions for both therapists and patients. The robot-assisted rehabilitation relies on repetitive therapy based on two lower-level interactive controllers depicted in Section IV-A. During each therapy iteration, the patient’s actual trajectories $\bm{x}_{t}$ are measured from the patient-side robot. The patient motion preferences are then encoded from recent actual trajectories by learning a probabilistic model (see Section IV-B). To minimize therapist assistance and encourage patient active participation, therapist-informed via-points are extracted from the measured corrective force, rather than transmitting the entire force profile or reference trajectory for patient-side robot control. The corresponding reference trajectory $\bm{x}_{r}$ for subsequent therapy iteration is then partially deformed according to the therapist-informed via-points, while preserving the patient’s motion preferences in the remaining segments (see Section IV-C). Finally, the therapist’s therapeutic skills are encoded as via-points in a latent space. A partial least squares regression function is fitted to reproduce these therapist-informed therapy skills (see Section IV-D).
@@ IV-A Lower-level Interactive Control
The robot dynamics in $n$ Degree of Freedom (DoF) Cartesian space are given by
@@
where $\bm{M}_{C,i}\in\mathbb{R}^{n\times n}$ denotes the inertia matrix of the robot, and $\bm{c}_{C,i}\in\mathbb{R}^{n}$ and $\bm{g}_{C,i}\in\mathbb{R}^{n}$ denote the Coriolis/centrifugal and the gravity vector, respectively. $\bm{f}_{cmp,i}\in\mathbb{R}^{n}$ is the control force to compensate the Coriolis/centrifugal and the gravity vectors. $\bm{f}_{ctrl,i}\in\mathbb{R}^{n}$ is the control command. Moreover, $\bm{f}_{ext,i}\in\mathbb{R}^{n}$ denotes the external force applied by human user to the robot. The subscript $_{i}$ is indexed by $i\in\{P,Th\}$, where $_{P}$ indicates the robot at the patient side, and $_{Th}$ for the therapist-side robot.
@@ IV-A1 Patient-side Robot Control
On the patient side, the robot aims to track the reference trajectory generated by the therapist input [15], thus the controller at the patient side is designed as,
@@ IV-A2 Therapist-side Robot Control
The robot on the therapist side aims to track the patient-side robot so that the therapist understand the situation of the patient [15]. Thus the controller on the therapist-side robot is designed as,
@@
where the $\bm{K}_{Th},\bm{D}_{Th}\in\mathbb{R}^{n\times n}$ are the positive definite stiffness and damping matrices for the therapist-side robot, respectively. And the interaction force between the therapist $\bm{f}_{ext,Th}$ will be the input of the trajectory generation.
@@ IV-B Patient Motion Preferences Encoding
For each therapy iteration, the therapist-side impedance controller in (3) is used to perform the movement therapy by tracking a reference trajectory $\bm{x}_{r}$ for $J$ times. Most recent actual trajectories of patient (e.g. gray lines in the Fig. 2(a)) are collected and down-sampled into a set of waypoints with length $N$ for learning a probabilistic model to encode patient motion preferences [31, 32]. The dataset $\mathcal{D}_{P}$ is constructed as $\mathcal{D}_{P}=\{\{\bm{\xi}_{n,j}^{i},\bm{\xi}_{n,j}^{o}\}_{n=1}^{N}\}_{j=1}^{J}$. $\bm{\xi}_{n,j}^{i}\in\mathbb{R}^{d_{i}}$ and $\bm{\xi}_{n,j}^{o}\in\mathbb{R}^{d_{o}}$ represent the input and output vectors with dimensions of $d_{i}$ and $d_{o}$, respectively. This study focuses on time-driven trajectories for movement therapy, where the input $\bm{\xi}_{n,j}^{i}$ corresponds to time $\bm{t}$, and the output $\bm{\xi}_{n,j}^{o}$ corresponds to the end-effector position $\bm{x}$ of the patient-side robot. GMMs are used to encode the mean and variance of motion preferences using collected waypoints in $\mathcal{D}^{i}_{P}$. The covariance reflects the variability of the actual trajectories [29]. The resulting joint probability distribution of time $\bm{t}$ and position $\bm{x}$ is given as $\mathcal{P}(\bm{t},\bm{x})=\sum\nolimits_{c=1}^{C}\varpi_{c}\mathcal{N}(\bm{t},\bm{x}|\bm{\mu}_{c},\bm{\Sigma}_{c})$. $C$ is the number of Gaussian components, which controls the smoothness of the fitted mean trajectory. The Gaussian component parameters $\{\bm{\mu}_{c},\bm{\Sigma}_{c},\varpi_{c}\}_{c=1}^{C}$ are characterized by its mean $\bm{\mu}_{c}$, covariance $\bm{\Sigma}_{c}$, and weight $\varpi_{c}$, estimated from $\mathcal{D}_{P}$ using the Expectation-Maximization algorithm (see green line and ellipses in Fig. 2(a)). The conditional probabilistic trajectory, denoted by $\{\widehat{\bm{x}}_{n}\}_{n=1}^{N}$, is then retrieved via GMR as $\mathcal{P}(\widehat{\bm{x}}_{n}|\bm{t}_{n})\sim\mathcal{N}(\widehat{\bm{\mu}}_{n},\widehat{\bm{\Sigma}}_{n})$. $\widehat{\bm{\mu}}_{n}$ and $\widehat{\bm{\Sigma}}_{n}$ are the conditional mean and covariance that can be calculated from the estimated GMMs parameters.
The parameterized potential trajectories of patient can be modeled as $\bm{x}(\bm{t})=\bm{\Theta}(\bm{t})^{T}\mathbf{w}$ [30]. $\bm{\Theta}\in\mathbb{R}^{d_{B}d_{o}\times d_{o}}$ is the basis feature matrix, $d_{B}$ is the dimensionality of the basis feature, and $\mathbf{w}\in\mathbb{R}^{d_{B}d_{o}}$ is the weight vector, assumed to follow a normal distribution $\mathbf{w}\sim\mathcal{N}(\bm{\mu}_{\mathbf{w}},\bm{\Sigma}_{\mathbf{w}})$ with unknown mean $\bm{\mu}_{\mathbf{w}}$ and covariance $\bm{\Sigma}_{\mathbf{w}}$. The parameterized probability of trajectory distribution is then given by
where $\mathcal{P}_{p}(\bm{x}|\bm{t}_{n})$ denotes the prior probability distribution of reference trajectory estimated from the dataset $\mathcal{D}_{P}$. Leveraging the properties of KL-divergence for Gaussian distributions, the mean and variance can be obtained by solving two sub-problems[30, 29].
IV-C Therapist-informed Trajectory Deformation
The therapist-side robot can reproduce the actual trajectory of patient for $J$ times during each therapy iteration. The patient motion preferences are estimated from the previous dataset $\mathcal{D}_{P}^{i}$ as in Section IV-B. For AAN therapy, the therapist applies corrective forces to partially deform the reference trajectory instead of updating the weight vector $\mathbf{w}$ to re-plan the entire reference trajectory. The therapist-informed corrective force $\bm{f}_{ext,Th}(\bm{t})$ is measured from the therapist-side robot during the reproduction of the patient’s actual trajectory. When $\bm{f}_{ext,Th}(\bm{t})$ exceeds a pretested threshold $|\overline{\bm{f}}|$, it indicates that the therapist intends to correct the therapeutic movement of the patient-side robot. As shown in Fig. 2(b), when a corrective force is activated at time $\bm{t}_{v}$, a therapist-informed via-point is inserted at $\bm{t}_{v}^{\prime}=\bm{t}_{v}+\delta\bm{t}$ to enable the partial deformation of reference trajectory. An additional dataset $\mathcal{D}_{T}$ that contains all therapist-informed via-points is introduced for each therapy episode to deform the reference trajectory for next therapy iteration. $\mathcal{D}_{T}$ shares the same structure as dataset $\mathcal{D}_{P}$ but consists of far fewer therapist-informed via-points ($L<<N$) to encourage the active participation of patients at other time. When the therapist-informed via-points $\{\bm{t}_{l},\bm{x}_{l}^{v}\}_{l=1}^{L}$ are assumed to follow a Gaussian distribution $\mathcal{P}_{v}(\bm{x}_{l}^{v}|\bm{t}_{l})\sim\mathcal{N}(\widehat{\bm{\mu}}_{l}^{v},\widehat{\bm{\Sigma}}_{l}^{v})$, the objective in (5) for estimating $\{\bm{\mu}_{\mathbf{w}},\bm{\Sigma}_{\mathbf{w}}\}$ can be reformulated following the idea in [30], as
where $\mathcal{P}_{p}(\bm{x}|\bm{t}_{n})$ and $\mathcal{P}_{v}(\bm{x}|\bm{t}_{l})$ denote the prior probability distribution estimated from datasets $\mathcal{D}_{P}$ and $\mathcal{D}_{T}$, respectively. To obtain a closed-form solution, as in (5), an extended dataset $\mathcal{D}_{U}=\mathcal{D}_{P}\cup\mathcal{D}_{T}$ is constructed, and the objective function is reformulated as
$\bm{\Sigma}(\bm{t}_{v}^{\prime})$ $=\bm{\Sigma}_{\mathbf{w}}(\bm{t}_{v}^{\prime})$
where $\beta_{\bm{\mu}}$ is a scaling factor determining the amplitude of the reference trajectory deformation. Furthermore, to enable smooth complete trajectory deformation, two additional via-points at $\bm{t}_{s}=0$ and $\bm{t}_{e}=\Delta T$ are inserted at the start and end points. $\Delta T$ is the duration of each therapy episode. The mean values of these via-points and the corresponding covariance matrices are defined as $\{\bm{x}_{e}(\bm{t}_{s}),\bm{x}_{e}(\bm{t}_{e})\}$ and $\{\bm{\Sigma}(\bm{t}_{s}),\bm{\Sigma}(\bm{t}_{e})\}$. The blue line in Fig. 2(b) represents the generated reference trajectory for the next therapy iteration, which preserves the patient’s previous motion preferences while adapting to the therapist-informed via-points.
IV-D Reproducing Therapist Input Skills
A latent space was constructed to represent the therapist’s input $\bm{\mathcal{\chi}}_{v}$ with the dimension of $M$. At each $i$-th therapy iteration, the therapist state is denoted as $\bm{s}_{i}$ with a dimension of $N$:
A dataset was constructed as $\mathcal{D}_{T}=\{(\bm{s}_{i},\bm{x}_{l,i}^{v})_{i=1}^{I}\}$. A partial least squares (PLS) regression function is utilized to reproduce the therapist input $\bm{Y}\in\mathbb{R}^{I\times M}$ from the collected feature $\bm{X}\in\mathbb{R}^{I\times N}$, as,
$\centering\small\bm{Y}=\bm{X}\mathbf{B}$ (10)
where $\mathbf{P}$ and $\mathbf{Q}$ are the loading matrices, $\mathbf{E}\in\mathbb{R}^{I\times N}$ and $\mathbf{F}\in\mathbb{R}^{I\times M}$ denote the residual matrices of $\bm{X}$and $\bm{Y}$, respectively. When only using several latent dimensions, the weight matrix is derived as
where $\mathbf{W}$ is the weight matrix of the input matrix.